OpenStudy (anonymous):

Find the derivative of f(x) = -6/x at x = 12

OpenStudy (bibby):

factor out the -6, so we're left with \(\dfrac{d}{dx}\dfrac{-6}{x}=-6\dfrac{d}{dx}\dfrac{1}{x}\) then we can rewrite \(\dfrac{d}{dx}\dfrac{1}{x}\implies \dfrac{d}{dx}x^{-1}\) and use the power rule

OpenStudy (anonymous):

so once I have d/dx x^-1 what do I plug in @bibby

OpenStudy (bibby):

-1 is now your n

OpenStudy (bibby):